OpenStudy (abbycross167):

Can someone please try to explain to me how to answer this question? Please Question~ Solve the systems using the elimination method. x + y = 0 3x + y = -4

OpenStudy (anonymous):

looks like a typo

OpenStudy (abbycross167):

What do you mean? what's wrong with it?

OpenStudy (anonymous):

take a look, you will see it

OpenStudy (misty1212):

Hi!!

OpenStudy (misty1212):

you got \[x+y=\] but nothing on the right hand side of the equal sign!

OpenStudy (abbycross167):

ooh I'm so sorry! I'll fix the question this is what should look like.... x + y = 0 3x + y = -4

OpenStudy (abbycross167):

@satellite73